OpenStudy (i0iz0gangster):

four questions and I have no idea what am doing 9. Evaluate the following expression when a = 2 and b = 5. (2 • a)b-1 10. Evaluate the following expression when a = 3 and b = -2. 7ab 11. If the mass of a proton is 2.53 x 10-37 gram, what is the mass of 1,000 protons? 12. What is the order of the mass of electrons from least to greatest? 12.4 x 10-3, 14 x 102, 0.013 x 103, and 157 x 10-4

OpenStudy (bahrom7893):

9) Plug in 2 and 5 instead of a and b: (2* 2)*5-1

OpenStudy (bahrom7893):

10) Again, plug in 3 instead of a, and -2 instead of b: 7*3*(-2)

OpenStudy (bahrom7893):

11) If one is 2.53 x 10-37 grams: 1 * p = p = 2.53 x 10-37 grams What is: 1,000 * p = 1,000 * 2.53 x 10-37 grams ?

OpenStudy (bahrom7893):

\[12.4 * 10^{-3}\]\[14*10^2\]\[0.013*10^3\]\[157*10^{-4}\] Which one of these numbers is the smallest, which one's the 2nd smallest, etc.?

OpenStudy (i0iz0gangster):

0.013 x 10^3 ?

OpenStudy (bahrom7893):

\[12.4*10^{-3}=0.0124\]\[14*10^2=1,400\]\[0.013*10^3=13\]\[157*10^{-4}=0.0157\]

OpenStudy (i0iz0gangster):

thank you soo much ur awesome and am sry for being a bit slow

OpenStudy (bahrom7893):

np

OpenStudy (i0iz0gangster):

and that's all the info that I need right

OpenStudy (bahrom7893):

yea, now just put those numbers in the right order, smallest to largest